# Parameter-Efficient Fine-Tuning with Attributed Patch Semantic Graph for Automated Patch Correctness Assessment

### v1

@@ Abstract
Automated program repair (APR) aims to automatically repair program errors without human intervention, and recent years have witnessed a growing interest on this research topic. While much progress has been made and techniques originating from different disciplines have been proposed, APR techniques generally suffer from the patch overfitting issue, i.e., the generated patches are not genuinely correct despite they pass the employed tests. To alleviate this issue, many research efforts have been devoted for automated patch correctness assessment (APCA). In particular, with the emergence of large language model (LLM) technology, researchers have employed LLM to assess the patch correctness and have obtained the state-of-the-art performance. The literature on APCA has demonstrated the importance of capturing patch semantic and explicitly considering certain code attributes in predicting patch correctness. However, existing LLM-based methods typically treat code as token sequences and ignore the inherent formal structure for code, making it difficult to capture the deep patch semantics. Moreover, these LLM-based methods also do not explicitly account for enough code attributes. To overcome these drawbacks, we in this paper design a novel patch graph representation named attributed patch semantic graph (APSG), which adequately captures the patch semantic and explicitly reflects important patch attributes. To effectively use graph information in APSG, we accordingly propose a new parameter-efficient fine-tuning (PEFT) method of LLMs named Graph-LoRA. Extensive evaluations have been conducted to evaluate our method, and the results show that compared to the state-of-the-art methods, our method improves accuracy and F1 score by 2.3% to 6.6% and 1.8% to 6.1% respectively.
@@ 1 Introduction
Software is unfortunately plagued with bugs, which can have serious consequences such as data loss, security flaw, and system hang. Resolving the software bugs is a notoriously difficult, expensive, and error-prone process [59, 77], and this issue is getting more and more serious as the scale and complexity of software continue to expand. To alleviate the burden on developers, the area of automated program repair (APR) arises and has received widespread attention from both academia and industry in the past two decades [57, 5]. The research agenda of APR is to automatically fix bugs in programs with less human intervention, and techniques originating from different disciplines have been proposed, remarkably including heuristic repair [35, 81, 65, 30], constraint-based repair [68, 41, 40], and learning-based repair [12, 75, 84, 79].
@@
Roughly speaking, APR techniques consist of three phases: fault localization [63, 76, 78], patch generation [31], and patch validation [10]. For the patch validation phase, the proposed techniques within the APR community typically evaluate the correctness of the generated patches using manually written test cases and a patch is deemed as correct in case it cam make the program pass all test cases. However, test cases in general can not fully specify the program behaviors and existing studies [17, 46, 39, 80, 50, 43, 21] demonstrate the existence of a significant portion of patches which pass the existing test suite but are actually incorrect. This phenomenon is called the patch overfitting problem, meaning that the generated patches simply overfit the existing test suite but do not achieve the expected program behavior in general.
@@
To alleviate this serious issue which overshadows the APR area, researchers have proposed many techniques for automated patch correctness assessment (APCA) [53, 73, 67, 50, 34, 52, 74]. Currently, APCA techniques can be broadly divided into two categories [17]: dynamic methods and static methods. Dynamic methods determine the correctness of patches by generating additional test cases and/or collecting features of test execution. For example, Opad [71] makes use of fuzz testing to generate new test cases and employs the corresponding oracles to enhance the patch correctness verification. Dynamic methods can achieve high accuracy but are very time-consuming due to the generation and/or running of tests. The static approach does not rely on running tests and instead assesses the patch correctness by the characteristics of the patch, such as its syntax and static semantic. For instance, on top of the assumption that the correct patch should be more similar to the defective code, S3 [33] measures the syntactic and semantic similarities between the patch and the error code to assess the patch correctness. Compared with dynamic approaches, static approaches can assess patch correctness quickly but suffer from the prediction accuracy issue. An ideal APCA approach should simultaneously maintain low time cost and high accuracy.
@@
Given that dynamic approaches necessarily take time to generate and/or run tests, much research attention has been devoted to static approaches in recent years and improving the accuracy of static approaches is viewed as a breakthrough [38, 83]. With the development of machine learning techniques, numerous learning-based APCA methods in the static category have been proposed in recent years. For instance, Ye et al. [73] manually design and extract 202 code features from the abstract syntax trees of defective code and patch code. Then, these code features and labels are given as inputs to three machine-learning models for constructing a probabilistic model. More recently, enlightened by the remarkable success of large language models (LLMs) for promoting code intelligence [26, 69], some researchers have employed LLMs to statically assess the correctness of patches. In particular, Zhou et al. [83] propose LLM4PatchCorrect, which predicts the patch correctness by feeding a LLM with information of labeled patches, such as error descriptions and failed tests. While these LLM-based methods have achieved state-of-the-art results in statically predicting patch correctness, drawbacks are associated with them. Previous works have demonstrated the importance of capturing patch semantic (including semantics of both the changed code and related unchanged code) in statically predicting patch correctness [62, 38], and an abundance of works also have shown that explicitly considering certain attributes associated with the changed and related unchanged code are extremely beneficial [73, 33, 64]. However, existing LLM-based methods typically treat code as token sequences and ignore the inherent formal structure for code, making it difficult to capture the deep patch semantics. Moreover, these LLM-based methods also do not explicitly account for enough attributes associated with the changed and unchanged code. Overall, these two drawbacks lead to the degraded performance in statically predicting patch correctness for LLM-based methods.
@@
To overcome the drawbacks, we in this paper propose a novel patch graph representation named Attributed Patch Semantic Graph (APSG). APSG is a directed graph which not only adequately captures the patch semantic through data and control flow between program elements, but also captures important attributes associated with the changed and related unchanged code by labeling different types of APSG nodes with different types of explicit attributes. Upon generating APSG, we further incorporate information of APSG into LLMs for statically predicting patch correctness. Inspired by the work of Yao et al. [72], we find that the attention mechanism can effectively merge graph information in APSG with sequence information in LLM. Besides, LLMs need to be fine-tuned in order to adapt to the APCA task. Taking these aspects into account, on top of LoRA [28]—one of the most advanced LLM parameter-efficient fine-tuning (PEFT) methods, we propose a new PEFT method called Graph-LoRA to retain graph information and fully train LLMs. Graph-LoRA can effectively fine-tune the parameters of LLMs and incorporate APSG information into LLMs through the attention mechanism.
@@
To verify the effectiveness of our method, we conduct experiments on four widely used datasets in the APCA field, including the Wang dataset [60], the Merge dataset [70], the Balance dataset [70], and the Lin dataset [38]. The results show that our method outperforms all static methods, including traditional methods and learning-based methods, in terms of accuracy, precision, recall, and F1 score. Compared with the best static method LLM4PatchCorrect [83], our method improves accuracy and F1 score by 2.3% to 6.6% and 1.8% to 6.1% respectively. In terms of precision, our method is closest to the dynamic method Opad [71]. The results of ablation studies show that both APSG and Graph-LoRA play a significant role in fine-tuning LLM on the APCA task. The results of cross-project prediction also show that our method achieves state-of-the-art performance when evaluating unseen patches.
@@ 2 Related Work
In recent years, automated patch correctness assessment (APCA) has gradually become a research hotspot in APR field. Depending on whether test case executions are needed, these APCA methods can be broadly divided into dynamic methods and static methods [17].
@@
Dynamic methods determine the correctness of patches by making use of test case generation techniques [4, 18] (particularly test amplification techniques that generate additional test cases [14]) and/or collecting features of test execution [49]. Yu et al. [80] give the definition of two different kinds of overfitting issues, that is, incomplete fixing and regression introduction, and the proposed dynamic methods typically have different strengths for them. Xin et al. [66] propose DiffTGen, which uses the test generation tool Evosuite [19] to generate additional test cases for enhancing the patch correctness check. PATCH-SIM, proposed by Xiong et al. [67] , uses a test generation tool to generate new test cases and assesses the correctness of the patch based on the similarity of the test case execution. The underlying assumption is that a correct patch should make the original program and the patched program behave similarly on the test cases that originally passed, but behave differently on the test cases that originally failed. Yang et al. [71] propose using fuzz testing to generate additional test cases and setting corresponding test assertions to validate the correctness of patches.
@@
In contrast, the static methods do not need to run test cases and the correctness of the patch is assessed by the characteristics of the patch. Le et al. [33] assume that the correct patch should be more similar to the defective code and propose S3 based on this assumption. On top of six features, S3 measures the syntactic and semantic similarities between the patch and the defective code to assess the patch correctness. Wen et al. [62] focus more on the contextual information of the patch and design three context-aware functions to assess the patch correctness. Xia et al. [64] first introduce entropy into the APCA task. They assume that the correct patches are more natural than the overfitting patches, and the entropy of patches can be used to measure patch correctness. With the development of machine learning techniques, many researchers have developed learning-based APCA methods. Ye et al. [73] manually design and extract 202 code features from the abstract syntax trees of the defective code and patch code. These code features and labels are then given as inputs to three machine-learning models for constructing a probabilistic model. Csuvik et al. [13] attempt to use BERT to generate embedding vectors to determine the similarity between the defective code and the patch code, thereby filtering overfitting patches. Zhang et al. [82] use the BERT [29] model as the encoder stack and use LSTM [25] to determine patch correctness. Tian et al. [54] use a neural network architecture to learn the semantic correlation between the bug reports and code patches to measure patch correctness. Tian et al. [52] additionally propose BATS, which predicts patch correctness based on the similarity of failed test cases. Lin et al. [38] use contextual and structural information to modify patch embeddings for improving the accuracy of patch correctness assessment.
@@
Most recently, enlightened by the remarkable success of large language models (LLMs) for promoting code intelligence, some researchers have employed LLMs to statically assess the correctness of patches. Notably, Zhou et al. [83] predict the patch correctness by feeding a LLM with information of labeled patches, such as error descriptions and failed tests.
@@
With the continuous development of deep learning technology and computing power, researchers have proposed various LLMs. In 2022, Google released a LLM called Chat-GPT [9], which demonstrates outstanding performance in the question-answering field. Touvron et al. [56] train a LLM called LLama using only public data, and release the model parameters to the research community. LLama has become the most popular open-source LLM. Roziere et al. [47] propose an open-source LLM for code, which has significantly improved performance for numerous tasks, including content filling, context information extraction, and instruction tracking. Li et al. [36] propose another open-source LLM for code named StarCoder, which expands the model input length to 8K and demonstrates excellent performance in the Python language.
@@
As LLMs become more common, it is particularly important to optimize computational efficiency and resource usage. The purpose of parameter-efficient fine-tuning (PEFT) is to reduce resource consumption during fine-tuning by training only a part of the model’s parameters. Houlsby et al. [27] add an adapter module to each layer of the pre-trained model, froze the main parameters of the model during fine-tuning and only fine-tune the newly added adapter structure. Inspired by the concept of prompt, Li et al. [37] propose prefix tuning, another fine-tuning method based on adding parameters. The method constructs a continuous and task-related prefix and only modifies the prefix of a specific task during model training. Guo et al. [23] propose a parameter-modifying fine-tuning method called Diffpruning, which describes fine-tuning as learning a diff vector and adding it to the pre-trained fixed model parameters. Hu et al. [28] assume that the model parameters can be updated by modifying the intrinsic rank of the parameters, and further propose an intrinsic rank adapter LoRA for fine-tuning LLMs. Based on LoRA, Chen et al. [11] propose LongLoRA, which splits the long context and processes each group of context separately through the shifted sparse attention mechanism.
@@ 3 Attributed Patch Semantic Graph
With the development of deep learning techniques, a significant portion of research efforts have been devoted to learning-based code intelligence and impressive results have indeed been obtained. One key to the success of these learning-based methods lies in appropriate code representation [2]. Currently, there are three main types of code representation methods within the literature: token-based method [1, 24], syntax-based method [3, 44], and semantic-based method [6, 7, 8, 22]. Token-based methods represent the code as a series of tokens, and this simple representation facilitates learning but limited semantics can be captured due to the ignorance of the inherent code structure. Syntax-based methods represent code in the form of trees, which contain rich semantic information but usually have a deep hierarchical structure. As a result, in practice, notable refinement efforts of the raw tree representation are typically required to make the learning a success. Semantic-based methods represent code in the form of graphs, which can effectively facilitate the capture of code semantics for learning models. Among the variety of proposed graph representations, notable ones include data flow graph [8], control flow graph [15], program dependence graph [7], and contextual flow graph [6].
In the APCA field, previous works have demonstrated the importance of capturing patch semantic (including semantics of both the changed code and related unchanged code) in statically predicting patch correctness [62, 38], and an abundance of works also have shown that explicitly considering certain attributes associated with the changed and related unchanged code are extremely beneficial [73, 33, 64]. While existing graph-based representations can effectively facilitate the code semantic learning, they typically do not simultaneously contain changed and unchanged code. In addition, existing graph-based representations do not involve any explicit code attributes. In light of these shortcomings, we in this paper design a novel directed patch graph representation named Attributed Patch Semantic Graph (APSG). For ease of reading, Fig. 1 gives an example of an Attributed Patch Semantic Graph for a simple patch.
Definition (Attributed Patch Semantic Graph) The Attributed Patch Semantic Graph for a patch is a triple tuple $<V,E,X>$ where $V$ is a set of nodes, $E$ is a set of directed edges between nodes in $V$, and $X$ is a mapping from nodes in $V$ to their attributes.
@@
Second, a directed edge in $E$ can be of 3 kinds: data flow edge, control flow edge, and sub-graph merge edge. The data flow and control flow edges established between statement nodes (i.e., nodes from the sets $V_{p}$, $V_{c}$, and $V_{ct}$) are similar to that of the typical program dependence graph. For control flow, there exists a control flow edge from node $a$ to node $b$ if $a$ represents the conditional statement whose predicate outcome directly controls whether $b$ is executed (the edge from node labeled with 4 to node labeled with 5 in Fig. 1 is an example). For data flow, there exists a data flow edge from node $a$ to node $b$ in case a certain variable $v$ defined at $a$ is used at $b$ and there is a path of the form $a$ $P$ $b$, where $P$ is a path along which $v$ is not redefined (the edge from node labeled with 2 to node labeled with 4 in Fig. 1 is an example, the involved variable is $x$). Previous studies [22, 55] have demonstrated the significance of considering data flow inside statements for accurately capturing code semantics, we thus also consider this aspect in APSG. In particular, there exists a data flow edge from node $a$ (in set $V_{v}$) to node $b$ (in set $V_{v}$) in case $a$ and $b$ correspond to variables on the right and left sides of an assignment statement respectively, and there exists a sub-graph merge edge from node $a$ (in set $V_{v}$) to node $b$ if $a$ corresponds to a variable on the left side of an assignment statement and $b$ corresponds to the assignment statement (the sub-graph for node labeled with 5 in Fig. 1 is an example).
@@
  • The patch node attribute is used to describe the characteristics of the patch at the line level, and we have considered four attributes for patch nodes. The first attribute is the edit distance, which calculates the number of times required for editing the defective code into the patch code based on the Manhattan distance. The second attribute is the patch entropy value, which describes the naturalness of a patch by calculating the maximum entropy of the patch and the calculation procedure follows that proposed by Xia et al. [64]. The third attribute is about the repair action, including addition, deletion, and replacement. The fourth property is the anti-pattern, which assesses whether the patch involves the forbidden transformations of the overfitting patches defined by Tan et al. [51].
@@
  • The variable node attribute represents the characteristics of the variable, and we have considered two attributes for variable nodes. The first attribute is variable type, which establishes the specified type for the variable, such as int, float, double, and bool. The second attribute is variable role, which describes the role of variables in computation and the calculation procedure follows that proposed by Du et al. [16]. As an example, in the code snippet “a + b”, the roles of variables a and b are MathOperatorLeft and MathOperatorRight, respectively.
@@
Based on the code analysis tool Spoon [42], we fully implement an analyzer to get APSG for a Java method and the analyzer supports modern Java versions up to Java 16.

### v2

@@ Abstract
Automated program repair (APR) aims to automatically repair program errors without human intervention, and recent years have witnessed a growing interest on this research topic. While much progress has been made and techniques originating from different disciplines have been proposed, APR techniques generally suffer from the patch overfitting issue, i.e., the generated patches are not genuinely correct despite they pass the employed tests. To alleviate this issue, many research efforts have been devoted for automated patch correctness assessment (APCA). In particular, with the emergence of large language model (LLM) technology, researchers have employed LLM to assess the patch correctness and have obtained the state-of-the-art performance. The literature on APCA has demonstrated the importance of capturing patch semantic and explicitly considering certain code attributes in predicting patch correctness. However, existing LLM-based methods typically treat code as token sequences and ignore the inherent formal structure for code, making it difficult to capture the deep patch semantics. Moreover, these LLM-based methods also do not explicitly account for enough code attributes. To overcome these drawbacks, we in this paper design a novel patch graph representation named attributed patch semantic graph (APSG), which adequately captures the patch semantic and explicitly reflects important patch attributes. To effectively use graph information in APSG, we accordingly propose a new parameter-efficient fine-tuning (PEFT) method of LLMs named Graph-LoRA. Our method focuses on the typical real-world scenarios where ground-truth patches are inaccessible, and does not rely on ground-truth patches to work. Extensive evaluations have been conducted to evaluate our method, and the results show that compared to the state-of-the-art methods, our method improves the accuracy and F1 score by 3.1% to 7.5% and 3.0% to 7.1% respectively.
@@ 1 Introduction
Software is unfortunately plagued with bugs, which can have serious consequences such as data loss, security flaw, and system hang. Resolving the software bugs is a notoriously difficult, expensive, and error-prone process [debugging, multiple-fault], and this issue is getting more and more serious as the scale and complexity of software continue to expand. To alleviate the burden on developers, the area of automated program repair (APR) arises and has received widespread attention from both academia and industry in the past two decades [urli2018design, 9393494]. The research agenda of APR is to automatically fix bugs in programs with less human intervention, and techniques originating from different disciplines have been proposed, remarkably including heuristic repair [le2011genprog, yuan2018arja, xin2019better, kim2013automatic], constraint-based repair [xuan2016nopol, nguyen2013semfix, mechtaev2015directfix], and learning-based repair [chen2019sequencer, ye2022neural, Recoder, yutse].
@@
Roughly speaking, APR techniques consist of three phases: fault localization [7390282, YUgui, 6113915], patch generation [6606626], and patch validation [9402121]. For the patch validation phase, the proposed techniques within the APR community typically evaluate the correctness of the generated patches using manually written test cases and a patch is deemed as correct in case it cam make the program pass all test cases. However, test cases in general can not fully specify the program behaviors and existing studies [overfittingsurvey, qi2015analysis, long2016analysis, yuemse, smith2015cure, 10.1145/3663529.3663776, 10.1145/3533767.3534368] demonstrate the existence of a significant portion of patches which pass the existing test suite but are actually incorrect. This phenomenon is called the patch overfitting problem, meaning that the generated patches simply overfit the existing test suite but do not achieve the expected program behavior in general.
@@
To alleviate this serious issue which overshadows the APR area, researchers have proposed many techniques for automated patch correctness assessment (APCA) [tian2020evaluating, ye2021automated, xiong2018identifying, smith2015cure, le2023invalidator, tian2022predicting, ye2021comprehensive]. Currently, APCA techniques can be broadly divided into two categories [overfittingsurvey]: dynamic methods and static methods. Dynamic methods determine the correctness of patches by generating additional test cases and/or collecting features of test execution. For example, Opad [yang2017better] makes use of fuzz testing to generate new test cases and employs the corresponding oracles to enhance the patch correctness verification. Dynamic methods can achieve high accuracy but are very time-consuming due to the generation and/or running of tests. The static approach does not rely on running tests and instead assesses the patch correctness by the characteristics of the patch, such as its syntax and static semantic. For instance, on top of the assumption that the correct patch should be more similar to the defective code, S3 [le2017s3] measures the syntactic and semantic similarities between the patch and the error code to assess the patch correctness. Compared with dynamic approaches, static approaches can assess patch correctness quickly but suffer from the prediction accuracy issue. An ideal APCA approach should simultaneously maintain low time cost and high accuracy.
@@
Given that dynamic approaches necessarily take time to generate and/or run tests, much research attention has been devoted to static approaches in recent years and improving the accuracy of static approaches is viewed as a breakthrough [lin2022context, zhou2024leveraging]. With the development of machine learning techniques, numerous learning-based APCA methods in the static category have been proposed in recent years. For instance, Ye et al. [ye2021automated] manually design and extract 202 code features from the abstract syntax trees of defective code and patch code. Then, these code features and labels are given as inputs to three machine-learning models for constructing a probabilistic model. More recently, enlightened by the remarkable success of large language models (LLMs) for promoting code intelligence [10.1145/3695988, 10713474], some researchers have employed LLMs to statically assess the correctness of patches. In particular, Zhou et al. [zhou2024leveraging] propose LLM4PatchCorrect, which predicts the patch correctness by feeding an LLM with information of labeled patches, such as error descriptions and failed tests. While these LLM-based methods have achieved state-of-the-art results in statically predicting patch correctness, drawbacks are associated with them. Previous works have demonstrated the importance of capturing patch semantic (including semantics of both the changed code and related unchanged code) in statically predicting patch correctness [wen2018context, lin2022context], and an abundance of works also have shown that explicitly considering certain attributes associated with the changed and related unchanged code are extremely beneficial [ye2021automated, le2017s3, xia2023automated]. However, existing LLM-based methods typically treat code as token sequences and ignore the inherent formal structure for code, making it difficult to capture the deep patch semantics. Moreover, these LLM-based methods also do not explicitly account for enough attributes associated with the changed and unchanged code. Overall, these two drawbacks lead to the degraded performance in statically predicting patch correctness for LLM-based methods.
@@
To overcome the drawbacks, we in this paper propose a novel patch graph representation named Attributed Patch Semantic Graph (APSG). APSG is a directed graph which not only adequately captures the patch semantic through data and control flow between program elements, but also captures important attributes associated with the changed and related unchanged code by labeling different types of APSG nodes with different types of explicit attributes. Upon generating APSG, we further incorporate information of APSG into LLMs for statically predicting patch correctness. Inspired by the work of Yao et al. [yao2020multimodal], we find that the attention mechanism can effectively merge graph information in APSG with sequence information in LLM. Besides, LLMs need to be fine-tuned in order to adapt to the APCA task. Taking these aspects into account, on top of LoRA [hu2021lora]—one of the most advanced LLM parameter-efficient fine-tuning (PEFT) methods, we propose a new PEFT method called Graph-LoRA to retain graph information and fully train LLMs. Graph-LoRA can effectively fine-tune the parameters of LLMs and incorporate APSG information into LLMs through the attention mechanism. Our method focuses on the typical real-world scenarios where ground-truth patches are inaccessible, and does not rely on ground-truth patches to work.
@@
To verify the effectiveness of our method, we conduct experiments on five APCA datasets, including the Wang dataset [wang2020automated], the Merge dataset [yang2023large], the Balance dataset [yang2023large], the Lin dataset [lin2022context], and the Multi-Benchmarks dataset. The first four datasets are based on the Defects4J benchmark [just2014defects4j], and have been widely used in the APCA field. The Multi-Benchmarks dataset is a large dataset we built that involves three bug benchmarks: Defects4J, Bugs.jar [saha2018bugs], and Bears [madeiral2019bears]. The experimental results show that our method outperforms all static methods, including traditional methods and learning-based methods, in terms of accuracy, precision, recall, and F1 score. Compared with the best static method LLM4PatchCorrect [zhou2024leveraging], our method improves the accuracy and F1 score by 3.1% to 7.5% and 3.0% to 7.1% respectively. In terms of precision, our method is closest to the dynamic method Opad [yang2017better]. The results of ablation studies show that both APSG and Graph-LoRA play a significant role in fine-tuning LLM on the APCA task. The results of cross-project prediction also show that our method achieves state-of-the-art performance when evaluating unseen patches.
@@ 2 Related Work
In the area of automated program repair (APR), there are typically two approaches for assessing the correctness of generated patches. The first approach is manual annotation where the correctness of generated patches is determined manually, and the second approach is automated assessment where manual efforts are not required. The study by Le et al. [le2019reliability] shows that manual annotation is more effective compared to automated assessment, but involves significant costs. Consequently, recent research efforts have been devoted primarily to automated patch correctness assessment (APCA), aiming to improve the effectiveness of APCA methods. Depending on whether test case executions are needed, these APCA methods can be broadly divided into dynamic methods and static methods [overfittingsurvey].
@@
Dynamic methods determine the correctness of patches by making use of test case generation techniques [47977, 6319229] (particularly test amplification techniques that generate additional test cases [danglot2017emerging]) and/or collecting features of test execution [shamshiri2015automatically]. Yu et al. [yuemse] give the definition of two different kinds of overfitting issues, that is, incomplete fixing and regression introduction, and the proposed dynamic methods typically have different strengths for them. Xin et al. [xin2017identifying] propose DiffTGen, which uses the test generation tool Evosuite [10.1145/2025113.2025179] to generate additional test cases for enhancing the patch correctness check. PATCH-SIM, proposed by Xiong et al. [xiong2018identifying] , uses a test generation tool to generate new test cases and assesses the correctness of the patch based on the similarity of the test case execution. The underlying assumption is that a correct patch should make the original program and the patched program behave similarly on the test cases that originally passed, but behave differently on the test cases that originally failed. Yang et al. [yang2017better] propose using fuzz testing to generate additional test cases and setting corresponding test assertions to validate the correctness of patches.
@@
In contrast, the static methods do not need to run test cases and the correctness of the patch is assessed by the characteristics of the patch. Le et al. [le2017s3] assume that the correct patch should be more similar to the defective code and propose S3 based on this assumption. On top of six features, S3 measures the syntactic and semantic similarities between the patch and the defective code to assess the patch correctness. Wen et al. [wen2018context] focus more on the contextual information of the patch and design three context-aware functions to assess the patch correctness. Xia et al. [xia2023automated] first introduce entropy into the APCA task. They assume that the correct patches are more natural than the overfitting patches, and the entropy of patches can be used to measure patch correctness. With the development of machine learning techniques, many researchers have developed learning-based APCA methods. Ye et al. [ye2021automated] manually design and extract 202 code features from the abstract syntax trees of the defective code and patch code. These code features and labels are then given as inputs to three machine-learning models for constructing a probabilistic model. Csuvik et al. [csuvik2020utilizing] attempt to use BERT to generate embedding vectors to determine the similarity between the defective code and the patch code, thereby filtering overfitting patches. Zhang et al. [zhang2024appt] use the BERT [kenton2019bert] model as the encoder stack and use LSTM [hochreiter1997long] to determine patch correctness. Tian et al. [tian2022change] use a neural network architecture to learn the semantic correlation between the bug reports and code patches to measure patch correctness. Tian et al. [tian2022predicting] additionally propose BATS, which predicts patch correctness based on the similarity of failed test cases. Lin et al. [lin2022context] use contextual and structural information to modify patch embeddings for improving the accuracy of patch correctness assessment.
@@
Most recently, enlightened by the remarkable success of large language models (LLMs) for promoting code intelligence, some researchers have employed LLMs to statically assess the correctness of patches. Notably, Zhou et al. [zhou2024leveraging] predict the patch correctness by feeding an LLM with information of labeled patches, such as error descriptions and failed tests.
@@
With the continuous development of deep learning technology and computing power, researchers have proposed various LLMs. In 2022, Google released an LLM called Chat-GPT [chatgpt], which demonstrates outstanding performance in the question-answering field. Touvron et al. [touvron2023llama] train an LLM called LLama using only public data, and release the model parameters to the research community. LLama has become the most popular open-source LLM. Roziere et al. [roziere2023code] propose an open-source LLM for code, which has significantly improved performance for numerous tasks, including content filling, context information extraction, and instruction tracking. Li et al. [li2023starcoder] propose another open-source LLM for code named StarCoder, which expands the model input length to 8K and demonstrates excellent performance in the Python language.
@@
As LLMs become more common, it is particularly important to optimize computational efficiency and resource usage. The purpose of parameter-efficient fine-tuning (PEFT) is to reduce resource consumption during fine-tuning by training only a part of the model’s parameters. Houlsby et al. [houlsby2019parameter] add an adapter module to each layer of the pre-trained model, froze the main parameters of the model during fine-tuning and only fine-tune the newly added adapter structure. Inspired by the concept of prompt, Li et al. [li2021prefix] propose prefix tuning, another fine-tuning method based on adding parameters. The method constructs a continuous and task-related prefix and only modifies the prefix of a specific task during model training. Guo et al. [guo2020parameter] propose a parameter-modifying fine-tuning method called Diffpruning, which describes fine-tuning as learning a diff vector and adding it to the pre-trained fixed model parameters. Hu et al. [hu2021lora] assume that the model parameters can be updated by modifying the intrinsic rank of the parameters, and further propose an intrinsic rank adapter LoRA for fine-tuning LLMs. Based on LoRA, Chen et al. [chen2023longlora] propose LongLoRA, which splits the long context and processes each group of context separately through the shifted sparse attention mechanism.
@@ 3 Attributed Patch Semantic Graph
With the development of deep learning techniques, a significant portion of research efforts have been devoted to learning-based code intelligence and impressive results have indeed been obtained. One key to the success of these learning-based methods lies in appropriate code representation [surveycode]. Currently, there are three main types of code representation methods within the literature: token-based method [Miltiadis2, Jinghui], syntax-based method [code2vec, Michael], and semantic-based method [Jakobovits, DependencyGraph, Kwonsoo, graphcodebert]. Token-based methods represent the code as a series of tokens, and this simple representation facilitates learning but limited semantics can be captured due to the ignorance of the inherent code structure. Syntax-based methods represent code in the form of trees, which contain rich semantic information but usually have a deep hierarchical structure. As a result, in practice, notable refinement efforts of the raw tree representation are typically required to make the learning a success. Semantic-based methods represent code in the form of graphs, which can effectively facilitate the capture of code semantics for learning models. Among the variety of proposed graph representations, notable ones include data flow graph [Kwonsoo], control flow graph [david2020neural], program dependence graph [DependencyGraph], and contextual flow graph [Jakobovits].
In the APCA field, previous works have demonstrated the importance of capturing patch semantic (including semantics of both the changed code and related unchanged code) in statically predicting patch correctness [wen2018context, lin2022context], and an abundance of works also have shown that explicitly considering certain attributes associated with the changed and related unchanged code are extremely beneficial [ye2021automated, le2017s3, xia2023automated]. While existing graph-based representations can effectively facilitate the code semantic learning, they typically do not simultaneously contain changed and unchanged code. In addition, existing graph-based representations do not involve any explicit code attributes. In light of these shortcomings, we in this paper design a novel directed patch graph representation named Attributed Patch Semantic Graph (APSG). For ease of reading, Fig. 1 gives an example of an Attributed Patch Semantic Graph for a simple patch.
Definition (Attributed Patch Semantic Graph) The Attributed Patch Semantic Graph for a patch is a triple tuple <V,E,X><V,E,X> where $V$ is a set of nodes, $E$ is a set of directed edges between nodes in $V$, and $X$ is a mapping from nodes in $V$ to their attributes.
@@
Second, a directed edge in $E$ can be of 3 kinds: data flow edge, control flow edge, and sub-graph merge edge. The data flow and control flow edges established between statement nodes (i.e., nodes from the sets $V_{p}$, $V_{c}$, and $V_{ct}$) are similar to that of the typical program dependence graph. For control flow, there exists a control flow edge from node $a$ to node $b$ if $a$ represents the conditional statement whose predicate outcome directly controls whether $b$ is executed (the edge from node labeled with 4 to node labeled with 5 in Fig. 1 is an example). For data flow, there exists a data flow edge from node $a$ to node $b$ in case a certain variable $v$ defined at $a$ is used at $b$ and there is a path of the form $a$ · $P$ · $b$, where $P$ is a path along which $v$ is not redefined (the edge from node labeled with 2 to node labeled with 4 in Fig. 1 is an example, the involved variable is $x$). Previous studies [graphcodebert, StructCoder] have demonstrated the significance of considering data flow inside statements for accurately capturing code semantics, we thus also consider this aspect in APSG. In particular, there exists a data flow edge from node $a$ (in set $V_{v}$) to node $b$ (in set $V_{v}$) in case $a$ and $b$ correspond to variables on the right and left sides of an assignment statement respectively, and there exists a sub-graph merge edge from node $a$ (in set $V_{v}$) to node $b$ if $a$ corresponds to a variable on the left side of an assignment statement and $b$ corresponds to the assignment statement (the sub-graph for node labeled with 5 in Fig. 1 is an example).
@@
  • The patch node attribute is used to describe the characteristics of the patch at the line level, and we have considered four attributes for patch nodes. The first attribute is the edit distance, which calculates the number of times required for editing the defective code into the patch code based on the Manhattan distance. The second attribute is the patch entropy value, which describes the naturalness of a patch by calculating the maximum entropy of the patch and the calculation procedure follows that proposed by Xia et al. [xia2023automated]. The third attribute is about the repair action, including addition, deletion, and replacement. The fourth property is the anti-pattern, which assesses whether the patch involves the forbidden transformations of the overfitting patches defined by Tan et al. [tan2016anti].
@@
  • The variable node attribute represents the characteristics of the variable, and we have considered two attributes for variable nodes. The first attribute is variable type, which establishes the specified type for the variable, such as int, float, double, and bool. The second attribute is variable role, which describes the role of variables in computation and the calculation procedure follows that proposed by Du et al. [yufse23]. As an example, in the code snippet “a + b”, the roles of variables a and b are MathOperatorLeft and MathOperatorRight, respectively.
@@
Based on the code analysis tool Spoon [spoon], we fully implement an analyzer to get APSG for a Java method and the analyzer supports modern Java versions up to Java 16.